OpenStudy (anonymous):

MEDAL AND FAN Can someone work through this with me and explain it please? The area of a rectangular piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. Part B: What is the width of the land in meters? Show the steps of your work.

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

Okay you would first have to set up 2 equations the equation for the area of a rectangle and an equation to set the length and width equal to each other given the variables. So your two equations would be \[l*w=280\] and \[l-5=w+1 \] *we know this because the sides of a square are equal so we would have to set the width and length equal to each other under the given conditions.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

and then we solve by substitution, so you have to isolate the variable. Since you are finding the width first let us isolate the length so we can use it to find the width

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

So the equation \[l * w =280\] becomes \[l=280/w\]

OpenStudy (anonymous):

but wouldn't I need to have another number in place of the w to get the length??

OpenStudy (anonymous):

no because , then we plug this into the other equation \[l-5=w+1\] and this becomes \[(280/w)-5 = w +1 \] Do I need to keep going?or do you have any more questions?

OpenStudy (anonymous):

would the (280 / w) - 5 = w + 1 be my equation for part A because this just does not make sense to me...

OpenStudy (anonymous):

Well you keep going and solve for w for the width. Here's a solution I found for you

OpenStudy (anonymous):

since it cant be -20 the answer should be 14 meters for part A

OpenStudy (anonymous):

Okay, that makes more sense. Thank you!!

OpenStudy (anonymous):

No problem, hope that helped :D do you still need help for part B are are you good?

OpenStudy (anonymous):

Part B would be 14 right? Because Part A just asks for the equation..

OpenStudy (anonymous):

Oh ya sorry just read that. Yup your good :D

OpenStudy (anonymous):

okay thanks!